OpenStudy (anonymous):

Identify the solutions of 2x2 + 13x + 6 = 0.

OpenStudy (anonymous):

use the quadratic equation \[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array}\] can you tell me what a b and c are?

OpenStudy (anonymous):

You can also try completing the square but I don't know if it's possible. Give that a try first and if you can't do it then use the quadratic equation.

OpenStudy (anonymous):

it's x = −1/2, x = −6